jennylove:

if im being truthful, I dont even know where to begin. So if someone could explain the steps ?

jennylove:

1 attachment
darkknight:

Do you know any of it?

darkknight:

So here is how the complex plane works drawing:  So I gave the location of 3-2i, so you should be able to figure out the quadrant

jennylove:

quadrant IV

darkknight:

Yes

darkknight:

yeah

jennylove:

then the last one would be quad ii ?

darkknight:

Because if you think about it this way: 3-2i(i^2) = 3i^2 - 2i^3 -3 - -i = -3+i

darkknight:

So that is why multiplying by i^2 would result in 180 degrees

darkknight:

So yes the last one is quadrant 2

jennylove:

ohh okay yes that makes sense! thank you for your help!

jennylove:

it seems like the answer was 90 degrees:/

jennylove:

could you explain why it could have been 90?

darkknight:

?

darkknight:

multiplying by i^2 is 90 degrees?

darkknight:

Oh no @jennylove you read the problem wrong. For the second one it was saying when a complex number is multiplied by an imaginary unit (ex i) then it would undergoe 90 degrees rotation

darkknight:

in this case because it is i^2 it is 180 degrees rotation

darkknight:

the second part was asking if it is multiplied by imaginary unit such as i not i^2

jennylove:

oh no!! I see the mistake!

darkknight:

Yeah, sorry I didn't catch that in the problem. I kinda read it as if it is multiplied by i^2 as in this case not an imaginary unit